OpenStudy (jkbo):

Please help.solve the system. 5x + 3y = 30 3x + 3y = 18 a.(6, -6) b.(0, -6) c.(-6, -6) d.(6, 0)

OpenStudy (amistre64):

let x=(9+n)/3 let y=(9-n)/3 plug in and solve for n

OpenStudy (jkbo):

thank you

OpenStudy (amistre64):

in the end, i believe (6,0) fits the best

OpenStudy (jkbo):

wait plug in 3 or 9? and thank you

OpenStudy (amistre64):

5x + 3y = 30 3x + 3y = 18 let x=(9+n)/3 let y=(9-n)/3 1: 5(9+n)/3 + 3(9-n)/3 = 30 2: 3(9+n)/3 + 3(9-n)/3 = 18 why did i pick those substitutions? 2: 9+n+ 9-n = 18, regardless of what n is so lets determine it from the other equation 1: 5(9+n)/3 + 3(9-n)/3 = 30 5(9+n) + 3(9-n) = 90 45 +5n + 27 -3n = 90 2n + 27 = 45 2n = 18 n = 9 x=(9+n)/3 =(9+9)/3 =18/3 =6 y=(9-n)/3 =(9-9)/3 =0/3 =0

OpenStudy (jkbo):

alright htank you so much

OpenStudy (amistre64):

youre welcome

OpenStudy (jkbo):

@amistre64 can you help me with this problem , Solve the following system by graphing and identify the point of intersection. -5x + 2y = -4 2x - y = 2 a.(–2, 0) b.(1, 2) c.(2, 1) d.(0, –2)